Circular No. 6011

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994U IN NGC 4948 Timothy Barker, Department of Physics and Astronomy, Wheaton

College, reports the discovery of a supernova during a CCD survey with a 0.36-m telescope, the object being located very close to the center of NGC 4948 (α = 13^{h}02^{m}.3, δ = −7^{o}41′, equinox 1950.0). When found on June 27.10 UT, SN 1994U was at mag ∼ 14; it was not visible on a similar image taken on June 10 (limiting mag ∼ 16.0) or on other images taken over the last 14 months.

S. Benetti, European Southern Observatory (ESO); and C. Zanin, Astronomy Department, Padua University, report that inspection of a fullyreduced CCD spectrogram (range 345-920 nm, resolution 0.6 nm), obtained on June 27.95 UT with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph), confirms that SN 1994U is a typical type-Ia supernova at maximum light. The spectrum is dominated by strong P-Cyg lines of intermediate-mass elements superimposed on a blue continuum. Visible lines include Ca II (H and K; infrared triplet); Si II (505.1, 597.2, 635.5 nm), Mg II (448.1 nm), S II (546.8, 561.2, 565.4 nm), and O I (777.3 nm). The expansion velocity deduced from the minimum of the Si II 635.5-nm line, corrected by the redshift (0.0048) of the parent galaxy and deduced by the Na I D interstellar lines, is ∼ 11200 km/s.

NOVA OPHIUCHI 1994 L. V. Morrison and R. W. Argyle, Royal Greenwich Observatory, report

the following accurate optical position for N Oph 1994, obtained from 14 observations with the Carlsberg Automatic Meridian Circle on La Palma: α = 17^{h}35^{m}44^{s}.681 ± 0^{s}.004, δ = −19^{o}19′34′′.02 ± 0′′.07 (equinox J2000.0, FK5 reference frame; epoch 1994.46). The following V magnitudes have also been obtained (± 0.05 mag): June 7.07 UT, 9.66; 9.07, 9.46; 10.07, 9.54; 11.06, 9.60; 12.06, 9.66; 13.06, 9.89; 16.05, 10.49; 17.05, 10.60; 18.04, 10.62; 19.04, 10.59; 20.04, 10.66; 21.03, 10.85; 25.02, 11.07; 26.02, 11.20.

NOVA SAGITTARII 1994 No. 2 Morrison and Argyle also report the following position for this nova

from 10 observations: α = 18^{h}30^{m}23^{s}.114 ± 0^{s}.004, δ = −17^{o}11′56′′.87 ± 0′′.08 (equinox J2000.0, FK5 reference frame; epoch 1994.46).

1994 June 28 Daniel W. E. Green